OpenStudy (anonymous):

1 explain in complete sentences which method you would use to solve the following system of equation 2 explain why you chose this method 3 provide the solution to the system X -3y+4z=-9 2x+y+3z=6 X-4y+3z=-9 Please help me

OpenStudy (triciaal):

To solve simultaneous equations with three variables it is best to eliminate one variable then solve with two variables by substitution. x -3y + 4z =-9 and x -4y + 3z =-9 x-3y +4z =x -4y +3z x-3y+4z-x+4y-3z = 0 y - z = 0 y = z substitute for y in eqn 1 x - 3y + 4y = -9 x + y =-9 substitute for y in eqn 3 x - 4y + 3y =-9 x-y =-9 add to get 2x = -18 x = -9 substitute for x to get y = 0 y = z = 0 sorry need to leave